mathslover (mathslover):

Find the ordered pairs of integer solutions (x,y) such that \(2^x + 1 = y^2\) .

mathslover (mathslover):

@vishweshshrimali5 @Miracrown

OpenStudy (vishweshshrimali5):

Couldn't find anything easier ? O.o

OpenStudy (vishweshshrimali5):

Okay one way is to draw the curves

mathslover (mathslover):

:D Well this question is quite straightforward. Just think about it! :-)

OpenStudy (vishweshshrimali5):

(3,3)

mathslover (mathslover):

No doubt in that but how did you get that?

OpenStudy (vishweshshrimali5):

(3,-3) Just trial and error method for now. :)

OpenStudy (vishweshshrimali5):

Actually not trial and error method

OpenStudy (vishweshshrimali5):

See: you have asked for integral solutions, so, 2^x + 1 must be an integer

OpenStudy (vishweshshrimali5):

That means that 2^x must also be an integer.

OpenStudy (vishweshshrimali5):

So, x must be +ve

OpenStudy (vishweshshrimali5):

Also, 2^x + 1 must be a square

mathslover (mathslover):

I'm getting your point, but that is not exactly the perfect solution for this. You can think other way too.

mathslover (mathslover):

May be, I will give miracrown and hartnn a chance to solve this. Till then, @vishweshshrimali5 - if you get any method, PM me. :-)

Miracrown (miracrown):

lol

Miracrown (miracrown):

you guys are brothers, why don't you do it one-on-one?

OpenStudy (vishweshshrimali5):

Actually, I have asked him this question again and again but he just prefers asking on OS !!

mathslover (mathslover):

Haha.. I know whenever I will ask to my brother, he will answer it immediately, so, I just discuss with him and others regarding the questions. :-)

mathslover (mathslover):

Have you given it a try @Miracrown ?

Miracrown (miracrown):

Well maybe he feels embarrassed asking you in a causal style.

OpenStudy (vishweshshrimali5):

May be its just that I don't like when he doesn't pay attention to me or tries to sleeps. Well it also may be because I scold him a lot while teaching him.

mathslover (mathslover):

Haha! I agree. I am sleepy .. lol and you scold me a lot- which is what I like .. OKay, we will discuss about this later, anyone knows the method ? ? ?

Miracrown (miracrown):

I do, but I won't help.

OpenStudy (vishweshshrimali5):

I really would have loved the graph method, but, I can't draw the graph with lappy.

OpenStudy (vishweshshrimali5):

*laptop

OpenStudy (vishweshshrimali5):

I am also not going to help you XD

mathslover (mathslover):

lol @Miracrown - Why so? Well, there is another method instead of graphing it.

mathslover (mathslover):

lol .. Okay, fine, I will put up my method here, as you guys *ARE NOT GOING TO HELP ME* : \(2^x + 1 = y^2 \\ 2^x = y^2 - 1\\ \text{Identity} \space a^2 - b^2 = (a+b)(a-b) \\ 2^x = (y+1)(y-1) \) Since the RHS has integers whose product is power of 2. Thus, both (y+1) and (y-1) should be the powers of 2. Now, as (y + 1 )- (y-1) = 2 therefore, it implies, y + 1 = 4 and y -1 = 2 this gives y = 3 and hence, x = 3. (3,3) is the solution.

Miracrown (miracrown):

Its really late, so excuse me if I go wrong anywhere.  the problem is asking for integer solutions.. let's analyses the x-exponent, it must be an integer let's try some values for x